# Unified Analysis of Screening Masses for Vector and Axial-Vector Mesons and Their Diquark Partners in the Contact Interaction Model

### v1

@@ IV SUMMARY AND PERSPECTIVE
  • Our results for the transverse screening masses of V and AV mesons were compared with those obtained using other approaches, showing good overall agreement. Nonetheless, although the CI framework predicts a smooth monotonic increase of the vector and axial-vector screening masses with temperature, several works in the literature report either decreasing or non-monotonic trends. Such discrepancies originate from the restricted validity of QCD sum rules, Refs. Veliev et al. (2011); Dominguez et al. (2013a), and effective chiral Lagrangians, Ref. Song (1993), which are only reliable for $T\leq T_{c}$ and rely on leading-order or hadronic-phase assumptions, thereby limiting their applicability near or beyond the transition point. Under contrast, the present CI treatment, applied uniformly to both light and heavy mesons, offers a consistent and complementary viewpoint that remains stable across the full temperature range considered.

### v2

@@ IV SUMMARY AND PERSPECTIVE
  • Our results for the transverse screening masses of V and AV mesons were compared with those obtained using other approaches, showing good overall agreement. Nonetheless, although the CI framework predicts a smooth monotonic increase of the vector and axial-vector screening masses with temperature, several works in the literature report either decreasing or non-monotonic trends. Such discrepancies originate from the restricted validity of QCD sum rules, Refs. Veliev et al. (2011); Dominguez et al. (2013b), and effective chiral Lagrangians, Ref. Song (1993), which are only reliable for $T\leq T_{c}$ and rely on leading-order or hadronic-phase assumptions, thereby limiting their applicability near or beyond the transition point. Under contrast, the present CI treatment, applied uniformly to both light and heavy mesons, offers a consistent and complementary viewpoint that remains stable across the full temperature range considered.